# HopaDIFF: Holistic-Partial Aware Fourier Conditioned Diffusion for Referring Human Action Segmentation in Multi-Person Scenarios

### v1

@@ 4.2 HopaDIFF
Diffusion-based Action Segmentation. ActDiff [31] first adopted diffusion to achieve the action segmentation, which aims to learn how to denoise the action segmentation ground truth iteratively perturbed by Gaussian noise during training time and generate concrete action segmentation labels based on random noise and temporal feature conditions during the inference time. To capture the underlying distribution of human actions, the model is trained to reconstruct the original action sequence from its corrupted form. During each training iteration, a random diffusion step $t\in\{1,2,...,T\}$ is selected. Subsequently, noise is added to the original ground truth sequence $\mathbf{y}_{0}$ based on a cumulative noise schedule, resulting in a corrupted version $\mathbf{y}_{t}\in\left[0,1\right]^{L\times C}$, which is defined as $\mathbf{y}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}}$ and $\epsilon$ denotes Gaussian noise. The denoising decoder $g_{\psi}$ is trained to handle sequences with varying noise levels, including fully random noise. At inference, it begins with a pure noise sample and progressively removes noise. The update at each step follows Eq. 2. This iterative process produces a sequence $\left[\hat{\mathbf{y}}_{T},\hat{\mathbf{y}}_{T-1},...,\hat{\mathbf{y}}_{0}\right]$, with $\hat{\mathbf{y}}_{0}$ as the final prediction approximating the ground truth. Though ActDiff shows promising performance on the RHAS task, the lack of target-aware partial feature reasoning and the low granularity of control for the action segmentation generation limit its performance on the RHAS task.
@@
To facilitate effective long-term information modeling and the exchange between holistic and partial cues related to the referred target, and to enhance the controllability of action segmentation generation conditioning, we propose a novel diffusion-based framework for referring human action segmentation, termed HopaDIFF. HopaDIFF is a novel diffusion-based framework for RHAS that integrates both holistic and partial video cues, guided by untrimmed video and corresponding textual references. It employs a dual-branch architecture, where the holistic branch captures global temporal context, and the partial branch focuses on fine-grained representations of the referred person using GroundingDINO [34] and VLM features from BLIP-2 [26]. To enhance temporal reasoning and controllability, HopaDIFF introduces a cross-input gate attentional module, i.e., HP-xLSTM, for holistic-spatial feature exchange, and incorporates a frequency-domain conditioning mechanism using DFT to guide the diffusion-based action segmentation process,
@@
Holistic and Partial Conditioned Diffusion for RHAS. HopaDIFF consists of two branches, i.e., holistic and partial reasoning branches. For the holistic reasoning branch, the original video, i.e., $\mathbf{v}_{h}$, is fed into a VLM-based cross-modal feature extractor (denoted as $\mathbf{F}_{\phi}$) together with textual reference, i.e., $\mathbf{r}$, to achieve the preliminary cross modal feature extraction. For the partial reasoning branch, the video and textual reference are first fed into GroundingDino (G-Dino) [34] to achieve the detection of the person of interest, and then the cropped frames based on the bounding boxes prediction are grouped and resized to formulate a new video, fed into the same VLM feature extractor later. Two encoders ($\mathbf{E}_{h}$, $\mathbf{E}_{p}$) with the same structure are used for these two branches, where the parameters are not shared. Then the cross-modal feature conditions can be achieved by Eq. 3.
@@
Holistic-Partial Aware xLSTM Temporal Reasoning. Extended LSTM (xLSTM) [4] introduces two major components, i.e., exponential gating and enhanced memory structures. These give rise to two variants, i.e., sLSTM, which uses scalar memory with mixing, and mLSTM, which employs matrix memory with a parallelizable outer-product update. While mLSTM drops recurrent memory mixing which can be used to enable parallelism, sLSTM supports multiple cells and heads with selective memory mixing across cells. The holistic partial feature exchange is conducted at the input gate of the mLSTM block, as Eq. 5.
@@
$\mathrm{i}_{t}^{m}$ $=\ \text{exp}\left(\tilde{\mathrm{i}}_{t}^{m}\right),\qquad\qquad\tilde{\mathrm{i}}_{t}^{h},~{}\tilde{\mathrm{i}}_{t}^{p}\ =\textbf{BCA}(\ \bm{w}^{h\top}_{\mathrm{i}}\ \bm{z}_{t}^{h}\ +\ b_{\mathrm{i}}^{h},\ \bm{w}^{p\top}_{\mathrm{i}}\ \bm{z}_{t}^{h}\ +\ b_{\mathrm{i}}^{h})$ input gate
@@
[…]
@@
Discrete Fourier Transformation based Frequency Conditions. Since action segmentation fundamentally relies on per-frame action recognition to produce temporally coherent segmentation outputs, a high degree of control granularity is essential when employing diffusion models to generate segmentation results from Gaussian noise, particularly when using features as conditions. Existing approaches, e.g., ActDiff [31], primarily leverage features in the temporal and spatial domains as conditioning inputs, while overlooking the potential advantages offered by the frequency domain. In this work, we explicitly incorporate frequency-domain information, derived from features aggregated via HP-xLSTM, to enhance the controllability of the diffusion process. The frequency-based conditions for both the holistic and partial branches are computed according to Eq. 6.
@@ 5.1 Implementation Details
All the experiments are conducted on $1$ NVIDIA A100 GPU and PyTorch 2.0.0. We adopt both BLIP-2 [26] and CLIP [43] as the feature extractors, for all the experiment tables shown in the main paper, we set frame length as $3,000$ for each sample to get the train/val/test sets, where cross-movie evaluation denotes that the videos used for the aforementioned three sets come from different movies, which measures both of the cross-subject generalizability and cross-scenario generalizability. On the other hand, random partition denotes that the samples of the train/val/test sets are randomly selected. For the cross-movie evaluation settings, the train/val/test sets consist of $566$, $61$, and $163$ long-video samples, respectively. For the random evaluation settings, the train/val/test sets consist of $626$, $82$, and $82$ long-video samples, respectively. Each sample has a time length of $3.3$ minutes. The input feature dimension of the encoder is $768$ for BLIP-2 [26] feature extractor and is $512$ for CLIP [43] feature extractor, where we use ASFormer ($12$ layers with $256$ feature maps) as the feature encoder and ASFormer ($8$ layers with $128$ feature maps) according to [57, 31]. All the models are trained using the Adam [8] optimizer with a batch size of $4$. The learning rate to train our model is selected as $41e-$ 4. The binary cross-entropy loss and the boundary loss are equally weighted. More details are delivered in the supplementary.
In line with prior studies in the field of human action segmentation [31, 37, 2], we evaluate model performance using frame-wise accuracy (ACC), edit score (EDIT), and segmental F1 scores at overlap thresholds {$10\%$, $25\%$, $50\%$} (F1@10, 25, 50). The ACC metric captures the accuracy of the predictions at the individual frame level, whereas the EDIT and F1 scores focus on the temporal consistency and quality of action segmentation between segments.
Table 2: Experimental results with BLIP-2 [26] cross-modal feature extractor and under random partition setting on the RHAS dataset.
@@
FACT [37] 26.30 0.27 55.86 54.06 50.38 26.08 0.27 52.91 50.77 47.06
ActDiff [31] 42.01 7.05 70.74 68.19 63.66 41.85 7.20 70.56 68.34 63.29
ASQuery [15] 31.37 0.10 35.21 33.90 31.94 32.30 0.12 35.59 33.58 29.51
LTContent [2] 35.72 0.30 68.70 66.40 62.25 34.23 0.31 64.70 63.09 58.50
RefAtomNet [39] 39.48 0.12 36.22 34.68 32.48 38.01 0.13 34.01 31.93 27.62
@@
Table 3: Experimental results with BLIP-2 [26] cross-modal feature extractor and under cross-movie evaluation setting on the RHAS dataset.
@@
FACT [37] 26.01 0.34 60.16 57.79 55.33 27.89 0.44 59.05 58.37 57.13
ActDiff [31] 4.46 5.96 38.86 38.36 37.51 2.36 15.09 22.44 22.28 21.80
ASQuery [15] 22.95 0.05 34.39 31.52 27.96 20.84 0.08 33.86 32.86 30.20
LTContent [2] 49.49 0.33 42.69 40.12 37.22 52.52 0.37 49.35 47.24 42.55
RefAtomNet [39] 35.00 0.10 49.58 46.60 41.58 38.73 0.13 41.16 39.17 35.44
@@
Table 4: Experiments with CLIP [43] and under random partition on RHAS133.
@@
FACT [37] 23.99 0.46 56.66 54.95 51.06 23.95 0.44 55.10 53.63 49.72
ActDiff [31] 13.31 7.28 29.88 28.84 25.92 13.21 7.38 29.19 28.19 25.68
ASQuery [15] 32.59 0.07 29.63 28.52 26.97 26.35 0.06 25.08 23.82 19.79
LTContent [2] 13.20 2.08 35.29 30.96 28.24 13.99 1.78 30.96 29.20 22.38
RefAtomNet [39] 27.27 0.10 29.38 28.25 25.92 27.80 0.11 29.27 27.86 24.41
@@
Table 5: Ablations on RHAS133 using BLIP-2 [26] and under partition setting on different movies.
@@ 5.2 Benchmark Analysis
We first conduct training and evaluations on the random partition of the RHAS dataset and use BLIP-2 [26] as the feature extractor, where the experimental results are shown in Table 2. The referring action recognition approach, i.e., RefAtomNet [39], shows the second-best ACC, i.e., $39.48\%$, while its performances of F1@{10, 25, 50} are limited, since it is not specifically designed for long video understanding and lacks of long temporal reasoning capability. Most of the human action segmentation baselines delivers better F1@{10, 25, 50}, where LTContent and ActDiff achieve $>60\%$ of F1@{10, 25, 50}, thanks to their promising long-term dependency modeling capability. Among all the human action segmentation baselines, ActDiff [31] shows best performance for the referring human action segmentation task, which delivers $42.01\%$ of ACC, $7.05\%$ of EDIT, $70.74\%$, $68.19\%$, abd $63.66\%$ of F1@{10, 25, 50} on the validation set and $41.85\%$ of ACC, $7.20\%$ of EDIT, $70.56\%$, $68.34\%$, and $63.29\%$ of F1@{10, 25, 50} on the test set.
Diffusion-based human action segmentation benefits from its diffusion-based framework, which iteratively refines action predictions from noise, enabling it to capture complex temporal dependencies and action priors, which benefits the cue reasoning of the referring human action segmentation task. However, we observe that limited performance is delivered by ActDiff [31] on the cross-movie evaluation setting, as shown in Table 3. One underlying reason is that its learned action priors, e.g., temporal position and relational dependencies, are subject- and scenario-specific and may not transfer effectively to movies with different scene structures, pacing, or narrative conventions. Among all the baselines used in the benchmark, LTContent [2] produces the best ACC as shown in Table 3, i.e., $49.49\%$, while FACT [37] delivers best performances in terms of F1@{10, 25, 50}, i.e., $60.16\%$, $57.79\%$, and $55.33\%$ on the validation set and $59.05\%$, $58.37\%$, and $57.13\%$ on the test set. Regarding the cross-movie evaluation, most of the benchmarked baselines suffer performance decays in terms of F1@{10, 25, 50}, indicating that cross-scenario and cross-subject setting is challenging for most of the human action segmentation baselines in the RHAS task, while the referring human action recognition baseline RefAtomNet [39] works better in cross-movie evaluation scenario, thanks to its cross-branch agent attention mechanism which could will address the information fusion between local and global cues. ActDiff delivers the best EDIT score on both the random partition and cross-movie partition due to its iterative denoising mechanism, which enhances temporal consistency and boundary precision across varying data distributions.
@@
We further evaluate the cross VLM backbone generalizability of all the leveraged baselines and our proposed HopaDIFF method in Table 4. We first observe that most of the leveraged approaches show better performances when they use the BLIP-2 [26] cross-modal feature extractor compared with the CLIP [43] cross-modal feature extractor, indicating that using a strong VLM backbone can benefit the RHAS task. Among all compared methods, our proposed HopaDIFF achieves the best performance across every metric on both validation and test sets. Specifically, HopaDIFF reaches $46.99\%$ ACC, $7.68\%$ EDIT, and $77.00\%$, $74.39\%$, $69.57\%$ F1 scores on the validation set, and $46.62\%$ ACC, $7.44\%$ EDIT, and $74.74\%$, $72.28\%$, $73.16\%$ F1 scores on the test set. These results surpass all baselines by large margins, particularly in temporal consistency and segmental precision. In contrast, the best baseline FACT only achieves $23.99\%$ ACC and a maximum F1@50 of $49.72\%$ on the test set, indicating the difficulty of using a weaker VLM backbone in our RHAS task. Notably, HopaDIFF maintains stable performance across diverse evaluation settings, highlighting its generalization capability. This performance gain is attributed to its dual-branch design that fuses holistic and partial cues, and the integration of Fourier-based frequency conditioning for finer generative control. Overall, the strong results confirm that HopaDIFF effectively addresses the unique challenges of referring human action segmentation in complex, multi-person scenarios.

### v2

@@ 4.2 HopaDIFF
Diffusion-based Action Segmentation. ActDiff liu2023diffusion first adopted diffusion to achieve the action segmentation, which aims to learn how to denoise the action segmentation ground truth iteratively perturbed by Gaussian noise during training time and generate concrete action segmentation labels based on random noise and temporal feature conditions during inference. To capture the underlying distribution of human actions, the model is trained to reconstruct the original action sequence from its corrupted form. During each training iteration, a random diffusion step $t\in\{1,2,...,T\}$ is selected. Subsequently, noise is added to the original ground truth sequence $\mathbf{y}_{0}$ based on a cumulative noise schedule, resulting in a corrupted version $\mathbf{y}_{t}\in\left[0,1\right]^{L\times C}$, which is defined as $\mathbf{y}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}}$ and $\epsilon$ denotes Gaussian noise. The denoising decoder $g_{\psi}$ is trained to handle sequences with varying noise levels, including fully random noise. In inference, it begins with a pure noise sample and progressively removes noise. The update at each step follows Eq. 2. This iterative process produces a sequence $\left[\hat{\mathbf{y}}_{T},\hat{\mathbf{y}}_{T-1},...,\hat{\mathbf{y}}_{0}\right]$, with $\hat{\mathbf{y}}_{0}$ as the final prediction approximating the ground truth. Though ActDiff shows promising performance on the RHAS task, the lack of target-aware partial feature reasoning and the low granularity of control for the action segmentation generation limit its performance on the RHAS task.
@@
To facilitate effective long-term information modeling and the exchange between holistic and partial cues related to the referred target, and to enhance the controllability of action segmentation generation conditioning, we propose a novel diffusion-based framework for referring human action segmentation, termed HopaDIFF. HopaDIFF is a novel diffusion-based framework for RHAS that integrates both holistic and partial video cues, guided by untrimmed video and corresponding textual references. It employs a dual-branch architecture, where the holistic branch captures global temporal context, and the partial branch focuses on fine-grained representations of the referred person using GroundingDINO liu2024grounding and VLM features from BLIP-2 li2023blip2 . To enhance temporal reasoning and controllability, HopaDIFF introduces a cross-input gate attentional module, i.e., HP-xLSTM, for holistic-spatial feature exchange, and incorporates a frequency-domain conditioning mechanism using DFT to guide the diffusion-based action segmentation process.
@@
Holistic and Partial Conditioned Diffusion for RHAS. HopaDIFF consists of two branches, i.e., holistic and partial reasoning branches. For the holistic reasoning branch, the original video, i.e., $\mathbf{v}_{h}$, is fed into a VLM-based cross-modal feature extractor (denoted as $\mathbf{F}_{\phi}$) together with textual reference, i.e., $\mathbf{r}$, to achieve the preliminary cross modal feature extraction. For the partial reasoning branch, the video and textual reference are first fed into GroundingDino (G-Dino) liu2024grounding to achieve the detection of the person of interest, and then the cropped frames based on the bounding boxes prediction are grouped and resized to formulate a new video, fed into the same VLM feature extractor later. Two encoders ($\mathbf{E}_{h}$, $\mathbf{E}_{p}$) with the same structure are used for these two branches, where the parameters are not shared. Then the cross-modal feature conditions can be achieved by Eq. 3.
@@
Holistic-Partial Aware xLSTM Temporal Reasoning. Extended LSTM (xLSTM) Beck2024xLSTMEL introduces two major components, i.e., exponential gating and enhanced memory structures. These give rise to two variants, i.e., sLSTM, which uses scalar memory with mixing, and mLSTM, which employs matrix memory with a parallelizable outer-product update. While mLSTM drops recurrent memory mixing which can be used to enable parallelism, sLSTM supports multiple cells and heads with selective memory mixing across cells. The holistic partial feature exchange is conducted at the input gate of the mLSTM block, as Eq. 5.
@@
$\mathrm{i}_{t}^{m}$ $=\ \text{exp}\left(\tilde{\mathrm{i}}_{t}^{m}\right),\qquad\qquad\tilde{\mathrm{i}}_{t}^{h},~\tilde{\mathrm{i}}_{t}^{p}\ =\textbf{BCA}(\ \bm{w}^{h\top}_{\mathrm{i}}\ \bm{z}_{t}^{h}\ +\ b_{\mathrm{i}}^{h},\ \bm{w}^{p\top}_{\mathrm{i}}\ \bm{z}_{t}^{h}\ +\ b_{\mathrm{i}}^{h})\ \ \$input gate
@@
$\mathbf{o}_{t}^{m}$ $=\ \sigma\left(\tilde{\mathbf{o}}_{t}^{m}\right)\ ,\qquad\qquad\quad\ \ \ \,\tilde{\mathbf{o}}_{t}^{m}\ =\ \bm{W}^{m}_{\mathbf{o}^{m}}\ \bm{z}_{t}^{m}\ +\ \bm{b}_{\mathbf{o}}^{m}$ $\text{output gate},$
@@
Discrete Fourier Transformation based Frequency Conditions. Since action segmentation fundamentally relies on per-frame action recognition to produce temporally coherent segmentation outputs, a high degree of control granularity is essential when employing diffusion models to generate segmentation results from Gaussian noise, particularly when using features as conditions. Existing approaches, e.g., ActDiff liu2023diffusion , primarily leverage features in the temporal and spatial domains as conditioning inputs, while overlooking the potential advantages offered by the frequency domain. In this work, we explicitly incorporate frequency-domain information, derived from features aggregated via HP-xLSTM, to enhance the fine-grained controllability of the diffusion process and the capability of the model to deal with complex periodic temporal patterns. The frequency-based conditions for both the holistic and partial branches are computed according to Eq. 6.
@@ 5.1 Implementation Details
All the experiments are conducted on $1$ NVIDIA A100 GPU and PyTorch 2.0.0. We adopt both BLIP-2 li2023blip2 and CLIP radford2021learning as the feature extractors, for all the experiment tables shown in the main paper, we set frame length as $3,000$ for each sample to get the train/val/test sets, where cross-movie evaluation denotes that the videos used for the aforementioned three sets come from different movies, which measures both of the cross-subject generalizability and cross-scenario generalizability. On the other hand, the random partition denotes that the samples of the train/val/test sets are randomly selected. For the cross-movie evaluation settings, the train/val/test sets consist of $566$, $61$, and $163$ long-video samples, respectively. For the random evaluation settings, the train/val/test sets consist of $626$, $82$, and $82$ long-video samples, respectively. Each sample has a time length of $3.3$ minutes. The input feature dimension of the encoder is $768$ for BLIP-2 li2023blip2 feature extractor and is $512$ for CLIP radford2021learning feature extractor, where we use ASFormer ($12$ layers with $256$ feature maps) as the feature encoder and ASFormer ($8$ layers with $128$ feature maps) according to 2021_BMVC_Yi ; liu2023diffusion . All the models are trained using the Adam diederik2014adam optimizer with a batch size of $4$. The learning rate to train our model is selected as 1e41e-4. The binary cross-entropy loss and the boundary loss are equally weighted. All the baselines in this paper adopt holistic input apart from the experiments reported in Table 6.
In line with prior studies in the field of human action segmentation liu2023diffusion ; lu2024fact ; bahrami2023much , we evaluate model performance using frame-wise accuracy (ACC), edit score (EDIT), and segmental F1 scores at overlap thresholds $10\%$, $25\%$, $50\%$} (F1@10, 25, 50). The ACC metric captures the accuracy of the predictions at the individual frame level, whereas the EDIT and F1 scores focus on the temporal consistency and quality of action segmentation between segments. Note that, our approach is a two-stage approach, where we rely on pre-extracted VLM features.
Table 2: Experimental results with BLIP-2 li2023blip2 cross-modal feature extractor and under random partition setting on the RHAS dataset.
@@
FACT lu2024fact 26.30 0.27 55.86 54.06 50.38 26.08 0.27 52.91 50.77 47.06
ActDiff liu2023diffusion 42.01 7.05 70.74 68.19 63.66 41.85 7.20 70.56 68.34 63.29
ASQuery gan2024asquery 31.37 0.10 35.21 33.90 31.94 32.30 0.12 35.59 33.58 29.51
LTContent bahrami2023much 35.72 0.30 68.70 66.40 62.25 34.23 0.31 64.70 63.09 58.50
RefAtomNet peng2024referring 39.48 0.12 36.22 34.68 32.48 38.01 0.13 34.01 31.93 27.62
@@
Table 3: Experimental results with BLIP-2 li2023blip2 cross-modal feature extractor and under cross-movie evaluation setting on the RHAS dataset.
@@
FACT lu2024fact 26.01 0.34 60.16 57.79 55.33 27.89 0.44 59.05 58.37 57.13
ActDiff liu2023diffusion 4.46 5.96 38.86 38.36 37.51 2.36 15.09 22.44 22.28 21.80
ASQuery gan2024asquery 22.95 0.05 34.39 31.52 27.96 20.84 0.08 33.86 32.86 30.20
LTContent bahrami2023much 49.49 0.33 42.69 40.12 37.22 52.52 0.37 49.35 47.24 42.55
RefAtomNet peng2024referring 35.00 0.10 49.58 46.60 41.58 38.73 0.13 41.16 39.17 35.44
@@
Table 4: Experiments with CLIP radford2021learning and under random partition on RHAS133.
@@
FACT lu2024fact 23.99 0.46 56.66 54.95 51.06 23.95 0.44 55.10 53.63 49.72
ActDiff liu2023diffusion 13.31 7.28 29.88 28.84 25.92 13.21 7.38 29.19 28.19 25.68
ASQuery gan2024asquery 32.59 0.07 29.63 28.52 26.97 26.35 0.06 25.08 23.82 19.79
LTContent bahrami2023much 13.20 2.08 35.29 30.96 28.24 13.99 1.78 30.96 29.20 22.38
RefAtomNet peng2024referring 27.27 0.10 29.38 28.25 25.92 27.80 0.11 29.27 27.86 24.41
@@
Table 5: Ablations on RHAS133 using BLIP-2 li2023blip2 and under partition setting on different movies.
@@
FACT lu2024fact 26.01 0.34 60.16 57.79 55.33 27.89 0.44 59.05 58.37 57.13
FACT lu2024fact w/ partial branch 18.35 0.32 62.57 61.10 61.10 18.34 0.43 63.72 63.35 62.40
ActDIFF liu2023diffusion 4.46 5.96 38.86 38.36 37.51 3.26 15.09 22.44 22.28 21.80
ActDIFF liu2023diffusion w/ DFT-cond 10.46 6.29 26.61 24.80 23.40 11.92 19.54 30.01 29.64 28.92
Cross-Attention Fusion chen2021crossvit 51.74 6.41 86.91 84.47 81.50 57.02 19.87 88.97 88.41 87.36
@@ 5.2 Benchmark Analysis
We first conduct training and evaluations on the random partition of the RHAS dataset and use BLIP-2 li2023blip2 as the feature extractor, where the experimental results are shown in Table 2. The referring action recognition approach, i.e., RefAtomNet peng2024referring , shows the second-best ACC, i.e., $39.48\%$, while its performances of F1@{10, 25, 50} are limited, since it is not specifically designed for long video understanding and lacks long temporal reasoning capability. Most of the human action segmentation baselines delivers better F1@{10, 25, 50}, where LTContent and ActDiff achieve $60\%$ of F1@{10, 25, 50}, thanks to their promising long-term dependency modeling capability. Among all the human action segmentation baselines, ActDiff liu2023diffusion shows best performance for the referring human action segmentation task, which delivers $42.01\%$ of ACC, $7.05\%$ of EDIT, $70.74\%$, $68.19\%$, abd $63.66\%$ of F1@{10, 25, 50} on the validation set and $41.85\%$ of ACC, $7.20\%$ of EDIT, $70.56\%$, $68.34\%$, and $63.29\%$ of F1@{10, 25, 50} on the test set.
Diffusion-based human action segmentation benefits from its diffusion-based framework, which iteratively refines action predictions from noise, enabling it to capture complex temporal dependencies and action priors, which benefits the cue reasoning of the referring human action segmentation task. However, we observe that limited performance is delivered by ActDiff liu2023diffusion on the cross-movie evaluation setting, as shown in Table 3. One underlying reason is that its learned action priors, e.g., temporal position and relational dependencies, are subject- and scenario-specific and may not transfer effectively to movies with different scene structures, pacing, or narrative conventions. Among all the baselines used in the benchmark, LTContent bahrami2023much produces the best ACC as shown in Table 3, i.e., $49.49\%$, while FACT lu2024fact delivers best performances in terms of F1@{10, 25, 50}, i.e., $60.16\%$, $57.79\%$, and $55.33\%$ on the validation set and $59.05\%$, $58.37\%$, and $57.13\%$ on the test set. Regarding the cross-movie evaluation, most of the benchmarked baselines suffer performance decays in terms of F1@{10, 25, 50}, indicating that cross-scenario and cross-subject setting is challenging for most of the human action segmentation baselines in the RHAS task, while the referring human action recognition baseline RefAtomNet peng2024referring works better in cross-movie evaluation scenario, thanks to its cross-branch agent attention mechanism which could will address the information fusion between local and global cues. ActDiff delivers the best EDIT score on both the random partition and cross-movie partition due to its iterative denoising mechanism, which enhances temporal consistency and boundary precision across varying data distributions. Since our task focuses on fine-grained action annotations, recognizing precise action boundaries is inherently difficult, leading to lower EDIT scores.
@@
We further evaluate the cross VLM backbone generalizability of all the leveraged baselines and our proposed HopaDIFF method in Table 4. We first observe that most of the leveraged approaches show better performances when they use the BLIP-2 li2023blip2 cross-modal feature extractor compared with the CLIP radford2021learning cross-modal feature extractor, indicating that using a strong VLM backbone can benefit the RHAS task. Among all compared methods, our proposed HopaDIFF achieves the best performance across every metric on both validation and test sets. Specifically, HopaDIFF reaches $46.99\%$ ACC, $7.68\%$ EDIT, and $77.00\%$, $74.39\%$, $69.57\%$ F1 scores on the validation set, and $46.62\%$ ACC, $7.44\%$ EDIT, and $74.74\%$, $72.28\%$, $73.16\%$ F1 scores on the test set. These results surpass all baselines by large margins, particularly in temporal consistency and segmental precision. In contrast, the best baseline FACT only achieves $23.99\%$ ACC and a maximum F1@50 of $49.72\%$ on the test set, indicating the difficulty of using a weaker VLM backbone in our RHAS task.
Notably, HopaDIFF maintains stable performance across diverse evaluation settings, highlighting its generalization capability. This performance gain is attributed to its dual-branch design that fuses holistic and partial cues, and the integration of Fourier-based frequency conditioning for finer generative control. Overall, the strong results confirm that HopaDIFF effectively addresses the unique challenges of referring human action segmentation in complex, multi-person scenarios. We further provided the number of trainable parameters, inference times, and GFLOPS of baselines and our HopaDIFF in Table 7 for efficiency comparison.